what is 30 to the 1.2 power
Answer by lwsshak3(6469)

what is 30 to the 1.2 power
30^1.2=30^(12/10)=30^(6/5)
=30 to the 6/5 power=5th root of 30 raised to the 6th power
using a calculator
5th root of 30=1.9744...
1.9744^6=59.2305...
Check:
30^1.2=59.2305...
It all depends what form you want it in.